# Erasing Without Remembering: Safeguarding Knowledge Forgetting in Large Language Models

### v1

@@ 6 Evaluation on UGBench
In this section, we evaluate PerMU on UGBench and analyze the results. The implementation parameters of PerMU are consistent with those of the baselines. We integrate GDR on retain set with $RW=1$. The noise ratio $P$ is set to 0.4, and we maintain $C=0.1$.
@@
6.1 Improved Unlearning Capabilities.
When tested on TOFU Forget05 dataset using LLaMA2-7B, as shown in Tab. II, PerMU demonstrates remarkable effectiveness in unlearning knowledge. Specifically, PerMU outperforms the baselines with a relative improvement of up to 50.13% (78.67 $\rightarrow$39.23) in Forget Probability and 14.43% (56.27$\rightarrow$64.39) on Forget Truth Ratio.
When tested on the Harry Potter dataset, as shown in Tab. VI, PerMU consistently outperforms baselines on both Phi-1.3B and LLaMA2-7B. Specifically, PerMU achieves an absolute improvement of up to 15.47% (87.35$\rightarrow$71.88) in Forget ROUGE, 8.69% (65.05$\rightarrow$56.36) in Forget Probability and 15.36% (84.90$\rightarrow$69.54) in Forget F1 when evaluated on LLaMA2-7B, all while maintaining high model utility. Furthermore, the performance improvements of PerMU are more pronounced on LLaMA2-7B compared to Phi-1.3B, highlighting its resilience to model scaling and its potential for application to larger models.
@@
TABLE VII: Experimental results for inverted relation data, subject-replaced data, and one-hop reasoned data on the ZsRE dataset using LLaMA2-7B. Most methods struggle to forget the “hard” in-scope knowledge that are logically related to the original unlearning sample. Nevertheless, PerMU achieves an improvement of up to 8.24% on FRT ratio (0.85$\rightarrow$0.92).
@@
PerMU 77.00 56.51 74.87 63.76 0.92 88.19 79.62 87.16 68.56 0.81 85.95 70.34 83.74 64.77 0.81
TABLE VIII: Although existing methods can effectively forget the original unlearning samples, they fail to generalise to forget rephrased unlearning samples (text in red). In contrast, PerMU demonstrates superior generalisation ability (text in green).
@@
6.2 Improved Generalisation Capabilities.
@@
When tested on the Rephrased TOFU dataset using LLaMA2-7B (Tab. IV), PerMU outperforms the baselines relatively by up to 43.53% (53.37$\rightarrow$30.14) in Forget Probability, and 20.81% (56.50$\rightarrow$68.26) in Truth Ratio. At the same time, PerMU preserves knowledge on the Real World dataset better than the baselines when using LLaMA2-7B, even surpassing the retain model by up to 10.25% (39.52$\rightarrow$49.77) on Real World Probability and 11.18%(52.76$\rightarrow$63.94) on Real World Truth Ratio (Tab. II). We present a case study of the generalisation issue. As shown in Tab. VIII, the predictions generated by ULD remain highly similar to the ground truth (highlighted in red), reflecting suboptimal generalization. In contrast, PerMU produces entirely different responses from the ground truth, demonstrating superior generalization ability.
@@
Furthermore, PerMU generalises effectively to rephrased Harry Potter data, achieving relative improvements of up to 17.18% (84.39$\rightarrow$69.89) in ROUGE and 19.87% (81.73$\rightarrow$65.49) in F1 when tested on LLaMA2-7B (Tab. VI), while incurring only a minor 1.34% (83.99$\rightarrow$82.65) reduction in model utility compared to the retain model.
@@
When evaluated on the ZsRE dataset, as shown in Tab. VII, although the ICL-based method demonstrates superior unlearning performance, it suffers from lower model utility, which limits its applicability in real-world scenarios. Apart from it, PerMU achieves the best FRT ratio across all three subsets, particularly excelling on inverted relation data, where it shows a relative improvement of up to 8.24% (0.85$\rightarrow$0.92) on FRT ratio. This highlights PerMU’s ability to strike a better balance between unlearning and retaining knowledge. However, the absolute generalisation performance of all methods in forgetting knowledge logically related to the original unlearning samples remains suboptimal, particularly when it comes to unlearning knowledge that requires one-hop reasoning, showing a gap of up to 2.23 times lower than the Retain model. Enhancing the absolute generalisation ability of the unlearned model is crucial in machine unlearning, leaving room for further exploration by the community.
6.3 Discussions.
We investigate whether PerMU effectively addresses the two underlying problems highlighted in Section 4. For the first problem, as shown in Fig. 6, the ranking of the first answer token in the next-token probability distribution of PerMU is significantly lower than that of other methods across layers, indicating that the unlearned model fails to recall the facts during inference, thereby addressing Cause 1. As for the second problem, PerMU produces a lower probability of generating rephrased answers on both unlearning samples and rephrased unlearning samples (as shown in Tab. III), with relative reductions of up to 4.1% and 9.4%, respectively. Moreover, the probability delta is 33.3% lower than that of other methods, effectively addressing Cause 2. These results confirm that PerMU exhibits a much more generalised unlearning capability.
@@
7 Comprehensive Study
In this section, we conduct a comprehensive study to evaluate the effectiveness of PerMU, which includes assessing the quality of the unlearned model’s generations. Additionally, we perform ablation studies to explore the impact of the perturbation ratio $P$, the tuning coefficient $C$, multiple subject scenarios, discrete-token-level perturbation, and the adoption of various retain losses.
7.1 Fluency Evaluation
@@
TABLE XI: The impact of the perturbation ratio $P$. Experiments are conducted on the TOFU Forget01 dataset using Phi-1.3B. Different values of $P$ lead to varying trade-offs, we select $P=0.4$ as the optimal perturbation ratio.
@@ 7.2 Perturbation Ratio $P$
As shown in Tab. XI, even a small amount of noise ($P=0.1$) is sufficient to achieve a notable unlearning effect. Moreover, increasing the noise ratio further enhances the effectiveness of unlearning. Specifically, when $P=1.0$, almost all metrics on the Forget and Rephrased Forget data achieve their best values. This is expected, as higher noise levels make it harder for the model to recall related facts, resulting in a more fact-unrelated probability distribution and better unlearning performance.
TABLE XII: The impact of the Tuning Coefficient $C$. Experiments are conducted on the TOFU Forget01 dataset using Phi-1.3B. Increasing $C$ results in an improved unlearning effect, but at the cost of decreased model utility. To balance model utility with effective unlearning, we select $C=0.1$ for our experiments.

### v2

@@ 6 Evaluation on UGBench
In this section, we evaluate PerMU on UGBench and analyze the results. The implementation parameters of PerMU are consistent with those of the baselines. We integrate GDR with $RW=1$, set the percentage of perturbed tokens to $K=0.4$, and maintain perturbation ratio $P=0.4$, coefficient $C=0.1$.
@@
6.1 Improved Unlearning Capabilities
When tested on TOFU Forget05 dataset using LLaMA2-7B, as shown in Tab. III, PerMU demonstrates remarkable effectiveness in unlearning knowledge. Specifically, PerMU outperforms the baselines with a relative improvement of up to 50.40% (78.67 $\rightarrow$39.02) in Forget Probability and 11.3% (56.27$\rightarrow$62.63) on Forget Truth Ratio.
When tested on the Harry Potter dataset, as shown in Tab. V, PerMU consistently outperforms baselines on both Phi-1.3B and LLaMA2-7B. Specifically, PerMU achieves an absolute improvement of up to 17.2% (87.35 $\rightarrow$ 70.15) in Forget ROUGE, 9.88% (65.05 $\rightarrow$ 55.17) in Forget Probability and 16.05% (84.90 $\rightarrow$ 68.85) in Forget F1 when evaluated on LLaMA2-7B, all while maintaining high model utility. Furthermore, the performance improvements of PerMU are more pronounced on LLaMA2-7B compared to Phi-1.3B, highlighting its resilience to model scaling and its potential for application to larger models.
@@ 6.1 Improved Unlearning Capabilities
TABLE VI: Experimental results for inverted relation data, subject-replaced data, and one-hop reasoned data on the ZsRE dataset using Phi-1.3B. Most methods struggle to forget the “hard” in-scope knowledge that are logically related to the original unlearning sample. Nevertheless, PerMU achieves an improvement of up to 9.41% on FRT ratio (0.85$\rightarrow$0.93).
@@
PerMU 73.95 55.34 72.52 62.54 0.93 86.53 75.29 85.51 67.13 0.81 85.63 68.05 83.68 63.83 0.81
$\textsc{PerMU}^{\dagger}$ 77.00 56.51 74.87 63.76 0.92 88.19 79.62 87.16 68.56 0.81 85.95 70.34 83.74 64.77 0.81
TABLE VII: Although existing methods can effectively forget the original unlearning samples, they fail to generalise to forget rephrased unlearning samples (text in red). In contrast, PerMU demonstrates superior generalisation ability (text in green).
@@
6.2 Improved Generalisation Capabilities
@@
When tested on the Rephrased TOFU dataset using LLaMA2-7B (Tab. IV), PerMU outperforms the baselines relatively by up to 40.73% (53.37 $\rightarrow$ 31.63) in Forget Probability, and 15.67% (60.75 $\rightarrow$ 70.27) in Truth Ratio. At the same time, PerMU preserves knowledge on the Real World dataset better than the baselines when using LLaMA2-7B, even surpassing the retain model by up to 10.7% (39.52 $\rightarrow$ 50.22) on Real World Probability and 6.84% (52.76 $\rightarrow$ 59.60) on Real World Truth Ratio (Tab. III). We present a case study of the generalisation issue. As shown in Tab. VII, the predictions generated by ULD remain highly similar to the ground truth (highlighted in red), reflecting suboptimal generalization. In contrast, PerMU produces entirely different responses from the ground truth, demonstrating superior generalization ability.
@@
Furthermore, PerMU generalises effectively to rephrased Harry Potter data, achieving relative improvements of up to 20.36% (84.39 $\rightarrow$ 67.20) in ROUGE and 22.72% (81.73 $\rightarrow$ 63.16) in F1 when tested on LLaMA2-7B (Tab. V), while incurring only a minor 1.5% (83.99 $\rightarrow$ 82.49) reduction in model utility compared to the retain model.
When evaluated on the ZsRE dataset, as shown in Tab. VI, although the ICL-based method demonstrates superior unlearning performance, it suffers from lower model utility, which limits its applicability in real-world scenarios. Apart from it, PerMU achieves the best FRT ratio across all three subsets, particularly excelling on inverted relation data, where it shows a relative improvement of up to 9.41% (0.85$\rightarrow$0.93) on FRT ratio. This highlights PerMU’s ability to strike a better balance between unlearning and retaining knowledge. However, the absolute generalisation performance of all methods in forgetting knowledge logically related to the original unlearning samples remains suboptimal, particularly when it comes to unlearning knowledge that requires one-hop reasoning, showing a gap of up to 2.23 times lower than the Retain model. Enhancing the absolute generalisation ability of the unlearned model is crucial in machine unlearning, leaving room for further exploration by the community.
Figure 6: Visualization of the model’s sensitivity to each token across all layers. Subject words are highlighted with brighter colors and exhibit higher MSM values, indicating the model’s increased sensitivity to these tokens.
6.3 Discussions
We investigate whether PerMU effectively addresses the two underlying problems highlighted in Section 4. For the first problem, as shown in Fig. 5, the ranking of the first answer token in the next-token probability distribution of PerMU is significantly lower than that of other methods across layers, indicating that the unlearned model fails to recall the facts during inference, thereby addressing Cause 1. As for the second problem, PerMU produces a lower probability of generating rephrased answers on both unlearning samples and rephrased unlearning samples (as shown in Tab. II), with relative reductions of up to 4.1% and 9.4%, respectively. Moreover, the probability delta is 33.3% lower than that of other methods, effectively addressing Cause 2. These results confirm that PerMU exhibits a much more generalised unlearning capability.
TABLE VIII: Comparison between the average MSM values of subject words and the other words on UGBench datasets. Subject words exhibit higher MSM values than other tokens, indicating greater sensitivity.
Dataset TOFU Harry ZsRE
Subject Words 0.000561 0.003219 0.005961
Other Words 0.000219 0.001708 0.003082
Ratio 2.66 1.86 1.93
6.4 Fast Alternative Implementation of PerMU
To reduce the computational overhead of identifying the top-$K$most sensitive tokens for each unlearning sample, we explore a more efficient implementation of PerMU. We begin by computing the MSM to determine the types of tokens to which the model is most sensitive. Experiments are conducted on UGBench datasets using Phi-1.3B.
As shown in Tab. VIII, the mean MSM value of the subject words is up to 2.66 times higher than that of other words. We further visualize this by normalizing the MSM values of all tokens in each sentence to a range between 0 and 1. As depicted in Fig. 6, the subject words consistently exhibit higher intensity values in the middle layers, indicating that the model is more sensitive to subject tokens. This observation aligns with the role of subject words in the mid-layer modules of generative language models, which are primarily responsible for recalling factual information [38, 49], thereby influencing the model’s output to reflect memorized knowledge. It also corresponds to the structure of multi-hop reasoning, which typically begins with subject enrichment [50]. These insights confirm that subject tokens are key carriers of factual memory.
Based on this analysis, we propose a fast variant of our method, $\textsc{PerMU}^{\dagger}$, which directly perturbs all subject tokens in the unlearning sample while keeping all other hyper-parameters unchanged. This approach not only effectively suppresses the model’s ability to recall facts related to subjects—particularly efficient in copyright and privacy-related datasets—but also significantly reduces training time. Specifically, for the TOFU and ZsRE datasets, where subject words are already labeled, we directly locate these tokens in the unlearning samples and apply perturbation. For the Harry Potter dataset, which contains numerous character names, we treat character names as subject words and perturb them accordingly. In contrast, for WMDP and MUSE, where subject words are either not apparent or are scattered across long-form sentences, we revert to the original PerMU method for unlearning.
We present the experimental results of $\textsc{PerMU}^{\dagger}$ on UGBench. As shown in Tab.IV, $\textsc{PerMU}^{\dagger}$ achieves unlearning performance comparable to existing methods, even surpasses PerMU by up to 17.64% (17.91 $\rightarrow$ 14.75) on Rephrased TOFU Forget01 Probability, and 8.19 (63.09 $\rightarrow$ 68.26) on Rephrased TOFU Forget05 TR. Although $\textsc{PerMU}^{\dagger}$ underperforms PerMU on Harry Potter dataset, it still outperforms baseline methods, achieving relative improvements of up to 15.47% (87.35 $\rightarrow$ 71.88) in Forget ROUGE, 8.69% (65.05 $\rightarrow$ 56.36) in Forget Probability and 15.36% (84.90 $\rightarrow$ 69.54) in Forget F1 when evaluated on LLaMA2-7B, all while maintaining high model utility. These results demonstrate the effectiveness of $\textsc{PerMU}^{\dagger}$ in achieving efficient forgetting with minimal computational overhead.
@@ 6.4 Fast Alternative Implementation of PerMU
7 Comprehensive Study
In this section, we conduct a comprehensive study to evaluate the effectiveness of PerMU. This includes assessing the quality of the unlearned model’s generations, evaluating performance on more diverse datasets and larger deep reasoning models, and analyzing computational overhead. Additionally, we perform ablation studies to investigate the impact of various hyper-parameters, including the percentage of perturbed tokens ($K$), the perturbation ratio ($P$), the tuning coefficient ($C$), discrete token-level perturbation, and the use of different retain losses.
7.1 Evaluation on Generation Quality
@@
TABLE XI: Experiment results on DeepSeek-R1-Distill-Llama-8B using the One-hop dataset. Deep reasoning models exhibit improved one-hop unlearning performance. However, the absolute gap between unlearning on the forget data and on the one-hop data remains substantial.
Dataset Forget Data Rephrased Forget Data One-hop Data MU
Metric RG ↓ Pr↓ F1↓ RG↓ Pr↓ F1↓ RG ↓ Pr↓ F1↓
GA+GD 57.65 38.93 53.31 59.87 40.30 55.88 75.23 58.75 72.50 59.69
DPO+GD 59.15 42.89 58.20 60.61 42.70 53.04 72.30 56.05 73.63 60.28
NPO+GD 57.52 35.19 52.87 58.91 39.89 53.07 74.23 54.95 72.32 59.85
TV 62.52 42.60 67.46 67.29 42.70 55.19 76.33 56.95 74.18 61.20
ULD 58.30 37.46 53.16 60.09 36.98 55.72 75.25 55.21 72.83 60.34
WHP 60.37 37.71 57.98 61.80 37.29 55.91 75.42 55.15 73.07 59.36
RMU 57.40 39.51 55.06 57.66 36.26 56.19 72.33 55.43 74.33 59.24
PerMU 56.21 34.95 52.38 54.95 34.58 53.10 73.82 55.30 72.58 58.78
TABLE XII: Evaluation on the WMDP and MUSE datasets using Vicuna-13B and MUSE-7B, respectively. PerMU consistently achieves the highest FRT score, with improvements of up to 10.77% (from 0.65 to 0.72).
Dataset WMDP (Vicuna-13B) MUSE (MUSE-7B)
Metric RG↓ Pr↓ F1↓ MU↑ FRT↑ RG↓ Pr↓ F1 ↓ MU↑ FRT↑
GA+GD 82.57 35.19 76.84 55.36 0.85 82.10 20.70 76.96 38.13 0.64
DPO+GD 82.42 35.57 75.37 54.14 0.84 84.73 38.94 75.50 34.19 0.51
NPO+GD 81.63 33.02 75.52 55.28 0.87 81.84 22.94 74.63 38.00 0.64
TV 82.95 34.70 77.36 54.64 0.84 90.60 20.03 77.37 36.61 0.58
ULD 81.04 32.17 76.62 55.38 0.88 89.26 31.76 74.75 41.51 0.64
WHP 82.57 33.64 74.94 56.25 0.89 83.20 23.90 69.31 38.20 0.65
RMU 83.16 38.93 78.49 60.14 0.90 91.92 62.16 78.89 38.80 0.50
PerMU 80.24 31.83 73.71 56.35 0.91 78.60 21.48 62.39 38.77 0.72
7.2 Evaluation on Larger Deep Reasoning Models
It is interesting to observe that models struggle to forget latent one-hop reasoning data (as shown in Tab. VI). Given the growing popularity of deep reasoning models such as OpenAI’s o1 and DeepSeek-R1, it is worth exploring whether these reasoning-enhanced models can achieve better forgetting generalisation on one-hop questions related to the forget data. To investigate this, we conduct experiments on DeepSeek-R1-Distill-Llama-8B using the ZSRE One-hop dataset. The model is first fine-tuned using LoRA with rank = 8, after which we apply various unlearning methods, updating only the LoRA parameters. As shown in Tab. XI, we have the following key observations:
[…]
In conclusion, deep reasoning models demonstrate improved one-hop unlearning performance. However, a substantial gap remains between the unlearning of forget data and one-hop data. Moreover, larger models face challenges in preserving utility while achieving effective unlearning, highlighting the need for further investigation in this area.
TABLE XIII: The impact of the percentage of perturbed tokens $K$ on Phi-1.3B. To balance forgetting and retention, we set $K=0.4$ in our main experiments.
TOFU Forget01 Dataset
Split Forget Rephrased Forget Retain MU↑ FRT↑
Metric RG↓ Pr↓ TR↑ RG↓ Pr↓ TR↑ RG↑ Pr↑ TR↑
K=0.1 47.19 37.96 58.74 40.94 32.32 59.51 82.04 88.47 48.45 51.71 1.31
K=0.2 44.98 35.66 60.50 40.20 29.35 59.83 79.18 87.27 47.96 51.55 1.38
K=0.3 43.93 32.75 61.64 39.75 28.41 61.38 80.70 87.34 47.41 51.92 1.43
K=0.4 43.38 30.58 63.76 39.70 27.69 61.71 79.17 87.21 47.98 51.21 1.44
K=0.5 42.99 30.29 64.87 38.41 25.09 63.41 78.42 87.42 48.08 51.42 1.50
K=0.6 42.85 30.02 65.96 37.63 26.71 64.05 77.77 85.76 46.67 51.55 1.50
K=0.7 42.49 30.51 66.06 36.76 26.04 65.54 77.24 85.29 46.36 52.17 1.54
K=0.8 42.19 28.39 67.52 37.00 25.21 66.56 77.73 85.36 46.31 51.82 1.56
K=0.9 41.89 29.31 67.21 36.98 26.13 66.85 77.17 85.08 46.19 50.39 1.50
K=1.0 41.57 27.84 68.46 36.74 24.5 68.08 76.28 84.63 46.24 50.24 1.54
7.3 Evaluation on More Diverse Datasets
To further evaluate the unlearning effectiveness, we conduct experiments on more challenging and diverse datasets—WMDP [23] and MUSE [24]—where knowledge is often distributed across long-form documents. For WMDP, we select 50 samples from BIO split as the forget set and 350 samples as the retain set. The target model is Vicuna-13B, fine-tuned using LoRA with rank = 8. For MUSE, we conduct experiments on BOOK split using MUSE-books-7B as the target model. To improve testing efficiency, we randomly sample 40 instances from the verbmem-forget split for unlearning and use the remaining 60 instances for retention. The average input length is 990.3 tokens, with a maximum length of 1031 tokens.
As shown in Tab. XII, RMU achieves the highest utility but at the cost of effective unlearning. This may be attributed to RMU’s design, which focuses on unlearning with limited samples rather than leveraging the entire forget set—thereby limiting its overall unlearning performance. In contrast, PerMU achieves a stronger forgetting effect, with improvements of 1.23% on WMDP and 3.24% on MUSE, and consistently attains the best FRT ratio. These results highlight the effectiveness of PerMU on larger models and longer-context scenarios.
TABLE XIV: Computational overhead of existing unlearning methods.
Metric Train step time (seconds) Memory Usage (MiB)
GA+KL 22.9342 35056
DPO+KL 47.1367 48132
NPO+KL 31.5571 36182
TV 14.2505 20442
ULD 32.4175 31014
WHP 24.1673 20442
PerMU 155.7616 52384
$\textsc{PerMU}^{\dagger}$ 24.3362 39054
7.4 Computational Overhead
We evaluated the computational overhead of existing methods from two perspectives: training step time and memory usage. The experiments were conducted on the Phi-1.3B model using the TOFU Forget01 dataset, with a batch size of 4 for 5 epochs. We report the total training step time and the maximum memory usage. As shown in Tab. XIV, although PerMU requires more training time, most of the overhead comes from computing the MSM to identify the top-$K$sensitive tokens. Notably, this step can be performed offline prior to training, significantly reducing both time and memory consumption during actual training. Moreover, our proposed fast variant, $\textsc{PerMU}^{\dagger}$, introduces no additional cost in training step time or memory usage, confirming its computational efficiency.
7.5 Percentage of Perturbed Tokens $K$
In PerMU, we identify the top-$K$ most sensitive tokens and apply perturbations to them. The choice of $K$ directly influences the unlearning effect. We analyze the impact of varying $K$ from 0.1 to 1.0 in increments of 0.1, where each value represents $K$of the tokens to be perturbed, while keeping all other hyper-parameters fixed. As shown in Tab. XIII, increasing $K$ consistently enhances the unlearning ability, with Forget Rouge improved by up to 5.62% (47.19 $\rightarrow$ 41.57) on the TOFU dataset. However, this improvement comes at the cost of reduced model utility. This trade-off is expected, as perturbing more tokens makes it harder for the model to recover and correctly predict the unlearning samples. Although the Forget Retain Trade-off (FRT) score peaks when $K=0.8$ to $0.9$, the performance on retained knowledge drops significantly—up to 5.76% in Rouge (82.04 $\rightarrow$ 76.28) and 3.84% in Probability (88.47 $\rightarrow$ 84.63). Such degradation may negatively impact the model’s utility on preserved knowledge. To strike a balance between forgetting and retention, we set $K=0.4$ in our main experiments, which achieves effective unlearning while preserving most of the model’s capacity to retain relevant knowledge.
@@ 7.5 Percentage of Perturbed Tokens $K$
TABLE XV: The impact of the perturbation ratio $P$. Experiments are conducted on the TOFU Forget01 dataset using Phi-1.3B. Different values of $P$ lead to varying trade-offs, we select $P=0.4$ as the optimal perturbation ratio.
@@ 7.6 Perturbation Ratio $P$
As shown in Tab. XV, even a small amount of noise ($P=0.1$) is sufficient to achieve a notable unlearning effect. Moreover, increasing the noise ratio further enhances the effectiveness of unlearning. Specifically, when $P=1.0$, almost all metrics on the Forget and Rephrased Forget data achieve their best values. This is expected, as higher noise levels make it harder for the model to recall related facts, resulting in a more fact-unrelated probability distribution and better unlearning performance.
TABLE XVI: The impact of the Tuning Coefficient $C$. Experiments are conducted on the TOFU Forget01 dataset using Phi-1.3B. Increasing $C$ results in an improved unlearning effect, but at the cost of decreased model utility. To balance model utility with effective unlearning, we select $C=0.1$ for our experiments.